# Low-temperature transport in high-conductivity correlated metals: a density-functional plus dynamical mean-field study of cubic perovskites

### v1

@@ III.6 Utilizing Fermi-liquid scaling for extracting transport properties
Figure 7: Real-frequency self energies $\Sigma(\omega)$ obtained from Padé analytic continuation for (a) SrVO3 and (b) SrRuO3 at $T=116$ K plotted as $-\text{Im}\Sigma(\omega)/(\pi T)^{2}$ versus $|\omega|/T$. $\eta$ denotes the frequency offset when evaluating the Padé approximats ($\Sigma(\omega+i\eta)$). The first 400 Matsuara data points were used to construct the Padé approximats. Solid (dashed) lines denote $\mathrm{Im}\Sigma(\omega)$ for $\omega>0$ ($\omega<0$).
@@
On the real-frequency axis, the FL form of the self-energy is defined in Eq. (10). In principle, the coefficient $C$ can be taken directly from evaluating $\mathrm{Im}\Sigma(\omega=0)$. However, when using imaginary-time impurity solvers requiring analytic continuation, we will demonstrate that evaluating $\mathrm{Im}\Sigma(\omega=0)$ can be quite sensitive to the continuation. As mentioned above, we choose the Padé analytic continuation to continue the Matsubara data to the real-frequency axis, which is simpler than other popular methods, like the maximum entropy method [114, 115, 81, 88, 89, 92, 94], and the community standard for obtaining quality self-energy data at low frequencies [111, 81]. This method has a broadening parameter $\eta$, which is used to evaluate the Padé approximants at $\Sigma(\omega+i\eta)$. For moderately correlated materials, the small el-el scattering rate makes the AC results sensitive to this choice. Figure 7(a) demonstrates this for the case of SrVO3 (the behavior of SrMoO3 and PbMoO3 is analogous), where we plot $-\text{Im}\Sigma(\omega)/(\pi T)^{2}$ versus $|\omega|/T$. Decreasing $\eta$ continuously decreases the magnitude of $\mathrm{Im}\Sigma(\omega=0)$ to clearly unphysical levels. This not only makes extracting $C$ impossible, but is also problematic for using $\mathrm{Im}\Sigma(\omega)$ in the full Kubo formula Eq. (1), which is sensitive to the low-$\omega$ behavior of the self-energy. We can notice in Fig. 7(a) that, for $|\omega|/T\gtrsim 5$, there is a region where $\mathrm{Im}\Sigma(\omega)$ is insensitive to $\eta$. Since we know the $\omega$ scaling of the FL self-energy [Eq. (10)], we can use this region, dominated by the $\omega^{2}$ behavior, to fit $C$ and extrapolate to small $\omega$. The result of the fitting is given by the gray region in Fig. 7(a), where the finite width denotes our estimated uncertainty in the fit. With this approach, we can obtain a robust value of $C$, and determine the correct behavior of $\mathrm{Im}\Sigma(\omega)$ at low frequencies. For $T$ outside of the FL regime, we generally do not know the $\omega$ scaling of $\mathrm{Im}\Sigma(\omega)$ and cannot use this method. However, in such materials with stronger correlations, the magnitude of $\mathrm{Im}\Sigma(\omega=0)$ is usually significantly larger, as is the case for SrRuO3 [Fig. 4(d)]. We show in Fig. 7(b) that this results in a straightforward convergence of the low-$\omega$ behavior of the self-energy with $\eta$, so that the AC in this region is expected to be reliable.
@@
A few additional comments are in order about the extraction of $C$. First, we note that the issues with analyzing the Matsubara self-energy mirror those of the real-frequency self-energy. For imaginary frequencies, the problematic $C^{(0)}$ estimate for $C$ is analogous to extrapolating the Matsubara data to $i\omega=0^{+}$. In a sense, taking the analytically continued data at low frequency is like performing this same extrapolation, since there are no Matsubara data below $\pi T$. By contrast, the $C^{(2)}$ estimate utilizes the higher-frequency data to obtain $C$, which is more robust for imaginary frequencies as it is for real frequencies.
@@
$C$ (eV-1) $\Phi(\epsilon_{\text{F}})$ (eV$/\Omega$cm) $A$ ($10^{-5}\mu\Omega\text{cm}/\text{K}^{2}$)
@@ IV Electron-electron contributions to resistivity in cubic oxides
Finally, we comment on the results for SrRuO3 in comparison with experiment. In Fig. 9, we plot experimental resistivities for SrRuO3 thin films and single crystals [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], subtracting the residual resistivity (determined by $\rho$ of the lowest extracted $T$ point), compared to our DFT+DMFT results (using QMC). In addition to the paramagnetic results (reproduced from Fig. 8), we include calculations that allow for ferromagnetic order. The latter shows a kink in $\rho$ at 290 K, which is roughly our calculated $T_{c}$ (as discussed above, it is expected that single-site DMFT overestimates $T_{c}$ by roughly a factor of two). Once in the ordered phase, $\rho$ is substantially lower than the paramagnetic result, consistent with the trends seen experimentally.
@@ V Summary and Outlook
Using cubic ABO3 perovskites as examples, we demonstrated a DFT+DMFT methodology for accurately extracting the electron-electron contribution to the DC resistivity in high-conductivity, moderately correlated materials. As for any material with a complex Fermi surface, converged Brillouin-zone integration is crucial for numerical precision. We demonstrate that adaptive integration methods are a powerful systematic way to achieve such convergence of the transport function. The key challenge specific to high-conductivity materials is that the magnitude of the low-$T$ scattering rates are very small. This requires special care when such quantities are extracted from the DMFT self-energies. We show that both low-frequency extrapolations of the Matsubara self-energies and Padé analytic continuation do not give accurate transport properties. Instead, we advocate leveraging the fact that these materials have moderate correlations, thus relatively high FL coherence scales. The known low-order frequency and temperature scalings can be exploited to extract accurate scattering rates. These scalings also allow QMC calculations at relatively high $T$ to be used to infer transport properties at very low $T$. Comparing results from complementary impurity solvers, e.g., ones that work in real frequencies (like NRG) or imaginary frequencies (like QMC), provides a strategy for validating the consistency of the analysis.
@@
We expect that similar strategies of utilizing FL properties will be useful for calculating other transport properties, e.g., the Seebeck coefficient, Hall conductivity, magnetoresistance, and optical conductivity. Moreover, by ensuring that the numerical aspects of the calculation are under control, this work sets the stage for a careful analysis of the theory of transport in correlated materials. Specifically, the role of vertex corrections to the DC conductivity in a framework beyond single-site DMFT remains to be investigated in the context of real materials. Previous works at the model level have shown that vertex corrections are important for low-dimensional systems [118, 119, 120, 121] and the low-density limit [122, 123].
@@
Finally, we show for SrRuO3 in this work and for SrVO3 and SrMoO3 in our companion paper [33] that the methods described here provide results that are qualitatively and quantitatively consistent with low-$T$ resistivity experiments, although such comparisons are challenging due to the spread in reported values and the difficulty of measuring small $\rho$ at low $T$. In Ref. 33, we emphasize that such results are crucial for correct interpretation of $\rho(T)$ data. Specifically, materials like SrVO3 and SrMoO3 exhibit a $T^{2}$ resistivity up to high $T$ ($>100$ K); a quantitative comparison with our DFT+DMFT results indicates that this cannot arise from FL electron-electron scattering as is sometimes assumed, but must be a result of electron-phonon scattering (not considered here or in Ref. 33[34]. Moreover, the comparison between $\rho$ of SrVO3, SrMoO3, and PbMoO3 indicates that the superlative conductivity of SrMoO3 is not a result of, e.g., weaker electron-electron correlations, and must also have its origins in electron-phonon scattering. Thereby, the calculations enabled by our methodology provide significant insight into the physics of transport in high-conductivity materials.
@@ Acknowledgements.
We acknowledge useful discussions with Andy Millis, Jernej Mravlje, and Jennifer Coulter. C.E.D. and J.L.H. acknowledge support from the National Science Foundation under Grant No. DMR-2237674. F.B.K. acknowledges funding from the Ministry of Culture and Science of the German State of North Rhine-Westphalia (NRW Rückkehrprogramm). The Flatiron Institute is a division of the Simons Foundation.
@@ Appendix A Crystal data and density-functional theory
DFT calculations are performed using the Vienna ab-initio Simulation Package (VASP) [52, 53, 54] with the exchange-correlation functionals of Perdew, Burke, and Ernzerhof [124]. For these calculations, the following valence states are treated explicitly: Sr ($4s,5s,4p,4d$), Mo ($4s,5s,4p,4d,4f$), Ru ($4s,5s,4p,4d,4f$), V ($3s,4s,3p,3d,4f$), and O ($2s,2p,3d$). The projector-augmented wave (PAW) method [125] is used to treat the core electrons. All DFT calulations are performed in the $Pm\overline{3}m$ structure with $21\times 21\times 21$ $k$-point mesh and a cutoff of $550$ eV. The lattice constants read SrVO3: 3.86 Å, SrMoO3: 4.00 Å, PbMoO3: 4.02 Å, and SrRuO3: 3.87 Å , which were determined by optimizing the unit cell using VASP.
@@ Appendix B Constrained RPA
The screened Coulomb interaction is calculated using the implementation of the constrained random-phase approximation (cRPA) in VASP [126], which uses maximally localized Wannier functions (MLWF) [55] obtained from Wannier90 [56]. For this calculation we used a $7\times 7\times 7$ $k$-point grid in the $Pm\overline{3}m$ cubic structure with a cutoff of $500$ eV. The interaction parameters are extracted by averaging the full four index interaction tensor to parameterize the Hubbard-Kanamori Hamiltonian [74].
@@ Appendix C QMC (TRIQS/cthyb)
Figure 10: NRG self-energy at $T=116$ K for SrVO3 and SrMoO3, obtained with the symmetric improved estimator (sIE) from Ref. 112 compared to the previously used improved estimator (IE) from Ref. 132.
@@ Appendix D NRG
Figure 11: Scattering-rate analysis of the electronic self-energy for PbMoO3 in (a) real and (b) Matsubara frequencies from QMC and NRG. (a) $-\mathrm{Im}\Sigma(\omega)/(\pi T)^{2}$ collapses onto the FL form $C[1+(\omega/\pi T)^{2}]$ which is shown as a gray shaded region. (b) $C$ estimated from Eq. (17) as a function of $T$.

### v2

@@ III.6 Utilizing Fermi-liquid scaling for extracting transport properties
Figure 7: Real-frequency self energies $\Sigma(\omega)$ obtained from Padé analytic continuation for (a) SrVO3 and (b) SrRuO3 at $T=116$ K plotted as $\text{Im}\Sigma(\omega)/(\pi T)^{2}$ versus $|\omega|/T$. $\eta$ denotes the frequency offset when evaluating the Padé approximats ($\Sigma(\omega+i\eta)$). The first 400 Matsuara data points were used to construct the Padé approximats. Solid (dashed) lines denote $\mathrm{Im}\Sigma(\omega)$ for $\omega>0$ ($\omega<0$). Note $\eta$ is units of meV.
@@
On the real-frequency axis, the FL form of the self-energy is defined in Eq. (10). In principle, the coefficient $C$ can be taken directly from evaluating $\mathrm{Im}\Sigma(\omega=0)$. However, when using imaginary-time impurity solvers requiring analytic continuation, we will demonstrate that evaluating $\mathrm{Im}\Sigma(\omega=0)$ can be quite sensitive to the continuation. As mentioned above, we choose the Padé analytic continuation to continue the Matsubara data to the real-frequency axis, which is simpler than other popular methods, like the maximum entropy method [117, 118, 84, 91, 92, 95, 97], and the community standard for obtaining quality self-energy data at low frequencies [114, 84]. This method has a broadening parameter $\eta$, which is used to evaluate the Padé approximants at $\Sigma(\omega+i\eta)$. For moderately correlated materials, the small el-el scattering rate makes the AC results sensitive to this choice. Figure 7(a) demonstrates this for the case of SrVO3 (the behavior of SrMoO3 and PbMoO3 is analogous), where we plot $\text{Im}\Sigma(\omega)/(\pi T)^{2}$ versus $|\omega|/T$. Decreasing $\eta$ continuously decreases the magnitude of $\mathrm{Im}\Sigma(\omega=0)$ to clearly unphysical levels. This not only makes extracting $C$ impossible, but is also problematic for using $\mathrm{Im}\Sigma(\omega)$ in the full Kubo formula Eq. (1), which is sensitive to the low-$\omega$ behavior of the self-energy. We can notice in Fig. 7(a) that, for $|\omega|/T\gtrsim 5$, there is a region where $\mathrm{Im}\Sigma(\omega)$ is insensitive to $\eta$. Since we know the $\omega$ scaling of the FL self-energy [Eq. (10)], we can use this region, dominated by the $\omega^{2}$ behavior, to fit $C$ and extrapolate to small $\omega$. The result of the fitting is given by the gray region in Fig. 7(a), where the finite width denotes our estimated uncertainty in the fit. With this approach, we can obtain a robust value of $C$, and determine the correct behavior of $\mathrm{Im}\Sigma(\omega)$ at low frequencies. For $T$ outside of the FL regime, we generally do not know the $\omega$ scaling of $\mathrm{Im}\Sigma(\omega)$ and cannot use this method. However, in such materials with stronger correlations, the magnitude of $\mathrm{Im}\Sigma(\omega=0)$ is usually significantly larger, as is the case for SrRuO3 [Fig. 4(d)]. We show in Fig. 7(b) that this results in a straightforward convergence of the low-$\omega$ behavior of the self-energy with $\eta$, so the AC in this region is expected to be reliable.
@@
A few additional comments are in order about the extraction of $C$. First, we note that the issues with analyzing the Matsubara self-energy mirror those of the real-frequency self-energy. For imaginary frequencies, the problematic $C^{(0)}$ estimate for $C$ is analogous to extrapolating the Matsubara data to $i\omega=0^{+}$. In a sense, taking the analytically continued data at low frequency is like performing this same extrapolation, since there are no Matsubara data below $\pi T$. By contrast, the $C^{(2)}$ estimate utilizes the higher-frequency data to obtain $C$, which is more robust for imaginary frequencies than it is for real frequencies.
@@
[…]
@@ IV Electron-electron contributions to resistivity in cubic oxides
Finally, we comment on the results for SrRuO3 in comparison with experiment. In Fig. 9, we plot experimental resistivities for SrRuO3 thin films and single crystals [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], subtracting the residual resistivity (determined by $\rho$ of the lowest extracted $T$ point), compared to our DFT+DMFT results (using QMC). In addition to the paramagnetic results (reproduced from Fig. 8), we include calculations that allow for ferromagnetic order. The latter shows a kink in $\rho$ at 290 K, which is roughly our calculated $T_{c}$ (as discussed above, it is expected that single-site DMFT overestimates $T_{c}$ by roughly a factor of two due to the lack of nonlocal correlations). Once in the ordered phase, $\rho$ is substantially lower than the paramagnetic result, consistent with the trends seen experimentally.
@@
Instead, we believe that much of the discrepancy is due to the fact that our NRG calculations were performed in the paramagnetic phase, and SrRuO3 is ferromagnetic at low $T$. As discussed above (see Fig. 9), $\rho$ is significantly larger in the paramagnetic phase, and, indeed, our calculations yield a larger $A$ and a smaller $T_{\text{FL}}$. We can also compare to CaRuO3, which is paramagnetic down to the lowest $T$ measured. The $A$ coefficient of CaRuO3 is somewhat larger (0.10.10.20.2 $\mu\Omega\,$cm/K2) [120], which is consistent with the stronger correlations in that material [83]. We have not attempted computing the $A$ coefficient in the ferromagnetic phase of SrRuO3, which would be necessary to quantify the effect of magnetic structure versus, e.g., nonlocal correlations and vertex corrections; the reason is that reaching the required very low $T$ is challenging for QMC and lifting the SU(2) spin symmetry makes multi-orbital NRG computations less efficient and/or accurate.
@@ V Summary and Outlook
Using cubic ABO3 perovskites as examples, we demonstrated a DFT+DMFT methodology for accurately extracting the local electron-electron contribution to the dc resistivity in high-conductivity, moderately correlated materials. As for any material with a complex Fermi surface, converged Brillouin-zone integration is crucial for numerical precision. We demonstrate that adaptive integration methods are a powerful systematic way to achieve such convergence of the transport function. The key challenge specific to high-conductivity materials is that the magnitude of the low-$T$ scattering rates are very small. This requires special care when such quantities are extracted from the DMFT self-energies. We show that both low-frequency extrapolations of the Matsubara self-energies and Padé analytic continuation do not give accurate transport properties. Instead, we advocate leveraging the fact that these materials have moderate correlations, thus relatively high FL coherence scales. The known low-order frequency and temperature scalings can be exploited to extract accurate scattering rates. These scalings also allow QMC calculations at relatively high $T$ to be used to infer transport properties at very low $T$. Comparing results from complementary impurity solvers, e.g., ones that work in real frequencies (like NRG) or imaginary frequencies (like QMC), provides a strategy for validating the consistency of the analysis.
@@
We expect that similar strategies of utilizing FL properties will be useful for calculating other transport properties, e.g., the Seebeck coefficient, Hall conductivity, magnetoresistance, and optical conductivity. Moreover, by ensuring that the numerical aspects of the calculation are under control, this work sets the stage for a careful analysis of the theory of transport in correlated materials. Specifically, the role of nonlocal correlations and vertex corrections to the dc conductivity in a framework beyond single-site DMFT remains to be investigated in the context of real materials. Previous works at the model level have shown that vertex corrections are important for low-dimensional systems [121, 122, 123, 124] and the low-density limit [125, 126].
@@
Finally, we show for SrRuO3 that the methods described here provide results that are qualitatively consistent with experimental resistivity, although such a comparison is challenging due to the spread in reported values and the difficulty of measuring small $\rho$ at low $T$ (this same problem occurs for SrVO3 and SrMoO3, see Ref. 33). Moreover, the comparison between $\rho$ of SrVO3, SrMoO3, and PbMoO3 indicates that the superlative conductivity of SrMoO3 is not a result of, e.g., weaker electron-electron correlations, and must also have its origins in electron-phonon scattering. Thereby, the calculations enabled by our methodology provide significant insight into the transport properties of high-conductivity materials.
@@ Acknowledgements.
We acknowledge useful discussions with Andy Millis, Jernej Mravlje, and Jennifer Coulter. C.E.D. and J.L.H. acknowledge support from the National Science Foundation under Grant No. DMR-2237674. F.B.K. acknowledges funding from the Ministerium für Kultur und Wissenschaft des Landes Nordrhein-Westfalen (NRW-Rückkehrprogramm). The Flatiron Institute is a division of the Simons Foundation. The data that support the findings of this article are openly available [127].
@@ Appendix A Crystal data and density-functional theory
DFT calculations are performed using VASP [52, 53, 54] with the exchange-correlation functionals of Perdew et al. [128]. For these calculations, the following valence states are treated explicitly: Sr ($4s,5s,4p,4d$), Mo ($4s,5s,4p,4d,4f$), Ru ($4s,5s,4p,4d,4f$), V ($3s,4s,3p,3d,4f$), and O ($2s,2p,3d$). The projector-augmented wave method [129] is used to treat the core electrons. All DFT calulations are performed in the $Pm\overline{3}m$ structure with $21\times 21\times 21$ $k$-point mesh and a cutoff of $550$ eV. The lattice constants read SrVO3: 3.86 Å, SrMoO3: 4.00 Å, PbMoO3: 4.02 Å, and SrRuO3: 3.87 Å , which were determined by optimizing the unit cell using VASP.
@@ Appendix B Constrained RPA
The screened Coulomb interaction is calculated using the implementation of the cRPA in VASP [130], which uses MLWFs [55] obtained from Wannier90 [56]. For this calculation we used a $7\times 7\times 7$ $k$-point grid in the $Pm\overline{3}m$ cubic structure with a cutoff of $500$ eV. The interaction parameters are extracted by averaging the full four index interaction tensor to parametrize the Hubbard-Kanamori Hamiltonian [74].
Figure 10: NRG self-energy at $T=116$ K for SrVO3 and SrMoO3, obtained with the symmetric improved estimator (sIE) from Ref. 115 compared to the previously used improved estimator (IE) from Ref. 131.
@@ Appendix C QMC (TRIQS/cthyb)
Figure 11: Scattering-rate analysis of the electronic self-energy for PbMoO3 in (a) real and (b) Matsubara frequencies from QMC and NRG. (a) $\mathrm{Im}\Sigma(\omega)/(\pi T)^{2}$ collapses onto the FL form $C[1+(\omega/\pi T)^{2}]$ which is shown as a gray shaded region. (b) $C$ estimated from Eq. (17) as a function of $T$.